OpenStudy (anonymous):

t^2+6t/t-1 - 7/t-1 simplify ?

OpenStudy (anonymous):

(t²+6t)/(t-1) - 7/(t-1) = (t²+6t-7)/(t-1) = ((t-1)(t+7))/(t-1) = (t-7)

OpenStudy (anonymous):

thanks for the website

OpenStudy (anonymous):

and the answer is t+7

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

search only a 8th grader lol website yahoo answers